OpenStudy (anonymous):

simplify by removing factors of one p^2-9/(p+3)^2

OpenStudy (anonymous):

Like this? \[\frac{p^{2}-9}{(p+3)^{2}} = \frac{(p-3)(p+3)}{(p+3)(p+3)} = \frac{p-3}{p+3}\]